For each of the following systems, sketch the x- and y-nullclines and use

this information to determine the nature of the phase portrait. You may

assume that these systems are defined only for x, y ≥ 0.

(d) x = x(2 – y – 2x), y = y(3 – y – 4x)
Answer #1

ye y(3-y-40 2 m a-y-am = 2m-y-an Jm.9tY(3-Y-ym) = 39-04my f(41y) Let 2 Let 2l2-9-2m) 0 2-yan0 0n 20) 3-3-4m)-0 3-Y-4m On 2 0At ) 1 he Mathix is O-1 Rigen valu semi stabu POint is So Relibrium ATAt (0,0) on (0,0) 2-m n 3-27-4n 2 Hene fhe gen value 0f the Matrx s 8,3 (o,0 is uneble Henee ahe 12 Mari x is the eigen value